OpenStudy (anonymous):

What is the product of 4x – 3 and x + 8?

OpenStudy (shamil98):

use foil (4x-3)(x+8)

OpenStudy (anonymous):

help me when done shamil98

OpenStudy (anonymous):

help me

OpenStudy (anonymous):

Yes, use FOIL (4x - 3)(x + 8) First: 4x * x = x^2 Outer: 4x * 8 = 32x Inner: -3 * x = -3x Last: -3 * 8 = -24 Now add like-terms. x^2 + 32x + -3x + -24 Final answer: x^2 + 29x - 24